# A Survey on State-of-the-art Deep Learning Applications and Challenges

### v5

@@ 3.2 Deep Unsupervised Learning
Deep unsupervised models are trained with an unlabelled dataset. The learning process of these models discovering patterns, structures and representations within the data without relying on explicit labels or supervision. Instead, these models often learn by optimizing objective functions that capture the underlying data characteristics such as clustering, learning useful feature embeddings and reconstructing input data from compressed representations. Examples of deep unsupervised models are autoencoders, generative adversarial networks and restricted Boltzmann machines.
@@
Autoencoder is a generative neural network model that learns to encode the input data into a compressed representation and then reconstructs the original data from this representation. The layers that encode the input data is known as encoder while the layers that responsible for the reconstruction is referred to as the decoder as shown in Fig. 16. The encoded data (hidden layer) represents the abstract features of the input data also known as latent space or encoding. The decoder can be removed from the autoencoder, creating a standalone model that can be used for data compression and dimensionality reduction [59], [60]. The decoder can also be replaced with predictive layers for classification task [61]. The network architecture can be built using fully-connected and convolutional layers [62].
@@
Several autoencoder variants have been introduced to improve the autoencoder’s ability to capture better feature representation. Some introduced penalty terms to the loss function such as sparsity penalty (sparse autoencoder) [63] to encourage sparse representation and Jacobian Frobenius norm (contractive autoencoder) [64] to be less sensitive to small and insignificant variations in the input data while encoding the feature representation. Others trained the autoencoder to recover original data from corrupted data with noise [65]. An improved denoising autoencoder knowns marginalized denoising autoencoder has been proposed which marginalizes the noise by adding a term that is linked to the encoding layer [66]. Variational autoencoder is a variant of autoencoder that has similar architecture i.e. encoder, latent space and decoder. Despite the similarity, instead of learning a fixed encoding, variational autoencoder learns the probability distribution of the input data in the latent space [67]. The model can be used to generate data by sampling from the learned probability distribution. The network architecture can be built by stacking more than one fully-connected layer and convolutional layer.
Generative Adversarial Network (GAN) is another generative neural network model that is designed for generating data that adheres closely to the distribution of the original training set. The model consists of two different neural networks namely generator and discriminator as shown in Fig. 17. The generator learns to imitate the distribution of the training set given a noise vector, effectively outsmarting the discriminator. Simultaneously during the training, the discriminator is trained to differentiate between the real data from the training set and synthetic data generated by the generator [68]. This intricate dynamic between the networks drives an iterative learning process whereby the generator continually refines its ability to create synthetic data that closely resembles the real data while the discriminator enhances its ability to distinguish between authentic and fake data.
@@
The model can be extended by providing the labels to both generator and discriminator in which the model known as conditional GAN, capable of generating 1000 image classes [69]. Conditional GANs require a labelled dataset which might limit its application. InfoGAN is similar to conditional GAN, but the labels are substituted with latent codes, which allows the model to be trained in an unsupervised manner [70]. GANs often suffer from mode collapse whereby the model can only generate a single or small set of outputs. Wasserstein GAN improves the training by utilizing Wasserstein loss function which measures the difference between the real and synthesized data distribution [71]. ProGAN tackles the training instability of GAN by progressively growing the generator and discriminator. The idea is that the model is scaled up gradually, starting with the simplest form of the problem and little by little the problem’s complexity is increased as the training progresses [72]. StyleGAN leverages the progressive GAN’s approach and neural style transfer to improve quality of the generated data [73]. The model is characterized by the independent manipulation of both style and content, allowing it to generate diverse styles and high quality data.
4 Applications of Deep Learning
@@ 4.1 Computer Vision
Computer vision is an essential field in artificial intelligence. It is a field of study that focuses on enabling computers to acquire, analyze and interpret visual inputs to derive meaningful information. The visual inputs can take many forms such as digital images, sequence of images or video and point cloud, and the source of these inputs can be camera, LiDaR, medical scanning machine etc. Deep learning, specifically CNN models have been widely used in real-world computer vision applications including image classification, object detection and image segmentation. This section discusses more details about the recent advancements in deep learning models that have been achieved over the past few years.
@@ 4.1.1 Image Classification
Image classification is a fundamental task in computer vision which involves categorizing an image into one of predefined classes based on the visual content. The objective of image classification is to enable computers or machines to differentiate between objects within images, in a manner similar to how humans interpret visual information. Image classification is a crucial component in various applications such as robotics, manufacturing and healthcare. LeNet-5, introduced in 1998, is one of the earliest convolutional neural networks that was successfully trained to classify handwritten digits. The model underwent a series of improvements, including the use of tanh and average pooling which enhanced its ability to extract hierarchical features, ultimately improving overall performance. The model architecture comprises two convolutional layers, each with an average pooling layer, followed by two fully-connected layers, including the output layer [79]. Since then, numerous CNN models have been proposed based on LeNet-5 for image classification [80, 81] but the most significant one is AlexNet in 2012 which saw a transformative breakthrough in deep learning. AlexNet is considered the first CNN model with a large number of parameters that significantly improved the performance of image classification on a very large dataset (ImageNet). The model won first place in ILSVRC 2012, improving the test error from the previous year by almost 10% [46]. Numerous significant CNN models have been introduced in subsequent ILSVRC competitions including ZFNet, VGG16, GoogleNet, ResNet and ResNext. In general, the research focused on increasing the number of layers, addressing the problem of vanishing gradient and diminishing of feature reuse.
@@
Vision transformers (ViT) offers an alternative to convolutional neural networks that have long been the dominant architecture for image classification, by leveraging self-attention mechanisms for scalable global representation learning. Despite its effectiveness, ViT is sensitive to hyperparameter optimization and substandard performance on smaller datasets [88]. Furthermore, ViT lacks the ability to leverage local spatial features which is inherent in convolutional neural networks [89]. Therefore, several studies attempt to incorporate convolutional layers into ViT architecture to improve its performance and robustness. In particular, conformer is a network architecture with two branches: CNN branch and a transformer branch to extract local and global features respectively [90]. Both branches are connected by two “bridges” of $1\times 1$ convolution and up or down sampling operations, allowing the branches to share their features and enhance the feature representation. Both branches output predictions which are combined to produce the final prediction. A hybrid architecture, named MaxViT, combines convolutional networks and vision transformer to address the lack of scalability issues of self-attention mechanisms when the model is trained on large input size [91]. The improved vision transformer is composed of two modules whereby the first module attends local features in non-overlapping image patches and the global features are attended by processing a grid of sparse and uniform pixels. The transformer is stacked with a block of convolutional layers to extract local spatial features. The architecture of MaxViT is shown in Figure 18. Another study proposes a convolutional transformer network, introducing the depthwise convolutional block into the ViT [92]. This configuration allows the model to exploit the ability of convolutional networks to extract local spatial features while the ViT attends the extracted local features to focus on relevant information, enhancing the model’s ability to capture complex patterns and relationships. Specifically, instead of linear mapping, depth-wise convolutional mapping is used to generate query, key and value matrices.
@@ 4.1.2 Object Detection
Deep learning plays a major role in significantly advancing the state-of-the-art in object detection performance. Region-based CNN (R-CNN) is the first breakthrough in object detection that combines CNN with selective region proposals [93]. The region proposals are the candidate bounding boxes serving as the potential region of interests (objects) within the input image, and the CNN are used to extract features from the region proposals and classify the regions for object detection. An improved model, named Fast R-CNN introduces two prediction branches: object classification and bounding box regression which improves the overall performance of object detection [94]. However, R-CNN and Fast R-CNN models are computationally expensive and slow, thus practically infeasible for real-time applications. Addressing this issue, Fast R-CNN is integrated with a region proposal network, referred to as Faster R-CNN [95]. The region proposal network (RPN) is used to efficiently generate region proposals for object detection. RPN takes an input image and output a set of rectangle object proposals, each with a confidence score to indicate the likelihood of an object’s presence. To this end, RPN introduces the concept of anchor boxes whereby multiple bounding boxes of different aspect ratios are defined over the feature maps produced by the convolutional networks. These anchor boxes are then regressed over the feature maps to localize the objects, contributing to the improved speed and effectiveness of Faster R-CNN. The training of Faster R-CNN is divided into two stages. First the RPN is pre-trained to generate the region proposals and then, the Fast R-CNN is trained using the region proposals generated by the RPN for object detection. The backbone network responsible for extracting the features for Faster R-CNN is either ZFNet or VGG16.
@@
In two-stage object detectors, the region proposals are generated first, and then used for object detection. The two-stage process is computationally intensive and infeasible for real-time object detection applications. You Only Look Once or YOLO proposes a one-stage detection by directly predicting bounding boxes and object’s confidence score in a single forward pass through the neural network [96]. This single pass architecture significantly reduces the computational complexity, making YOLO suitable for real-time object detection applications. In YOLO, the input image is divided into S×S grids, each grid cell is responsible for detecting the objects present in the cell. Specifically, each grid cell predicts multiple bounding boxes and associated object’s confidence score, enabling simultaneous object detection across the entire image. Subsequent enhancements such as YOLOv3 [97] and YOLOv4 [98] are proposed, improving the model’s capability and accuracy. Single Shot Multibox Detector (SSD) is another one-stage detector, aims to address the issue of real-time object detection [99]. SSD also eliminates the region proposal generation and directly predicts bounding boxes and confidence scores, reducing the computational complexity. To improve the overall performance, SSD produces the predictions from different levels of feature maps, allowing detection of objects of different sizes in the input image.
@@
A common issue in object detection problems is the extremely imbalanced ratio of foreground to background classes. Addressing this issue, RetinaNet introduces a loss function that is based on the cross entropy called focal loss. Focal loss reduces the loss contribution of easily classified objects, allowing the model training to focus on the difficult objects [100]. RetinaNet adopts the Feature Pyramid Network (FPN) [101] with ResNet as the backbone network for extracting the feature maps. FPN is a network architecture with a pyramid structure that efficiently captures multiscale feature representation, facilitating object detection across various sizes. To further improve the overall performance, EfficientDet introduces bi-directional FPN which incorporates multi-level feature fusion to better capture multiscale feature representation [102]. Also, the model utilizes EfficientNet [103] as the backbone network to achieve a balance between computational efficiency and accuracy.
@@
Object detection performance often relies on a post-processing step called non-maximum suppression (NMS) to eliminate duplicate detections and select the most relevant bounding boxes. Specifically, NMS sorts all detection boxes based on their confidence scores, selects a box with the maximum score and discards the other boxes with a significant overlap with the selected box. This process is repeated on the remaining detection boxes. However, due to the inconsistency between the confidence score and the quality of object localization, NMS retains poorly localized bounding boxes with high confidence score while discarding more accurate predictions with poor confidence score. To mitigate this limitation, instead of discarding the neighboring boxes with significant overlap, soft-NMS applies Gaussian function to lower their confidence scores [104]. The idea is not to discard the neighboring bounding boxes, but gradually decline their scores based on the extend of the overlap with the selected box. This results in a smoother suppression, preserving the better-localized bounding boxes. Adaptive NMS introduces an adaptive threshold for the suppression of bounding boxes [105]. The algorithm dynamically adjusts the threshold based on the level of overlapping of the selected box with the other bounding boxes.
@@
Detection Transformer (DeTR) is an end-to-end trainable object detection model that leverages the transformer architecture to eliminates the need for handcrafted components such as anchor boxes and non-maximum suppression [106]. The self-attention mechanism of the transformer captures the global context and relationships between different parts of the image, allowing it to localize the objects and remove duplicate predictions. The model is trained with a set of loss functions that perform bipartite matching between the predicted and ground truth objects. DeTR uses ResNet as backbone network. Despite the success of DeTR in simplifying and improving object detection tasks, DeTR suffers from a long training time and low performance at detecting small objects due to its reliance on self-attention mechanism of the transformer, which lacks a multiscale feature representation. To mitigate this limitation, Deformable DeTR introduces a multiscale deformable attention module which can effectively capture feature representation at different scales [107]. Furthermore, the attention module leverages deformable convolution, allowing the model to adapt to spatial variation and capture more informative features in the input data. Dynamic DeTR addresses the same issues by utilizing a deformable convolution-based FPN to learn multiscale feature representation [108]. Moreover, the model replaces the transformer encoder with a convolution-based encoder to attend to various spatial features and channels. This modification allows the model to effectively detect small objects and converge faster during training. The architecture of dynamic DeTR is shown in Figure 19. A training scheme known as Teach-DeTR is proposed to improve the overall performance of DeTR [109]. The training scheme leverages the predicted bounding boxes by other object detection models during the training by calculating the loss of one-to-one matching between the object queries and the predicted boxes.

### v6

@@ 3.2 Deep Unsupervised Learning
Deep unsupervised models are trained with an unlabeled dataset. The learning process of these models discovering patterns, structures, and representations within the data without relying on explicit labels or supervision. Instead, these models often learn by optimizing objective functions that capture the underlying data characteristics such as clustering, learning useful feature embeddings and reconstructing input data from compressed representations. Examples of deep unsupervised models are autoencoders, generative adversarial networks and restricted Boltzmann machines.
@@
An autoencoder is a generative neural network model that learns to encode the input data into a compressed representation and then reconstructs the original data from this representation. The layers that encode the input data are known as encoder while the layers that responsible for the reconstruction are referred to as the decoder as shown in Fig. 20. The encoded data (hidden layer) represents the abstract features of the input data, also known as latent space or encoding. The decoder can be removed from the autoencoder, creating a standalone model that can be used for data compression and dimensionality reduction [76], [77]. The decoder can also be replaced with predictive layers for classification task [78]. The network architecture can be built using fully connected and convolutional layers [79].
@@
Several autoencoder variants have been introduced to improve the autoencoder’s ability to capture better feature representation. Some introduced penalty terms to the loss function such as sparsity penalty (sparse autoencoder) [80] to encourage sparse representation and Jacobian Frobenius norm (contractive autoencoder) [81] to be less sensitive to small and insignificant variations in the input data while encoding the feature representation. Others trained the autoencoder to recover original data from corrupted data with noise [82]. An improved denoising autoencoder knowns marginalized denoising autoencoder has been proposed which marginalizes the noise by adding a term that is linked to the encoding layer [83]. Variational autoencoder is a variant of autoencoder that has similar architecture: encoder, latent space and decoder. Despite the similarity, instead of learning a fixed encoding, variational autoencoder learns the probability distribution of the input data in the latent space [84]. The model can be used to generate data by sampling from the learned probability distribution. The network architecture can be built by stacking more than one fully connected layer and convolutional layer.
Generative Adversarial Network (GAN) is another generative neural network model that is designed for generating data that adheres closely to the distribution of the original training set. The model consists of two different neural networks namely generator and discriminator as shown in Fig. 21. The generator learns to imitate the distribution of the training set given a noise vector, effectively outsmarting the discriminator. Simultaneously, during the training, the discriminator is trained to differentiate between the real data from the training set and synthetic data generated by the generator [85]. This intricate dynamic between the networks drives an iterative learning process whereby the generator continually refines its ability to create synthetic data that closely resembles the real data, while the discriminator enhances its ability to distinguish between authentic and fake data.
@@
The model can be extended by providing the labels to both generator and discriminator in which the model known as conditional GAN, capable of generating 1000 image classes [86]. Conditional GANs require a labelled dataset, which might limit its application. InfoGAN is similar to conditional GAN, but the labels are substituted with latent codes, which allows the model to be trained in an unsupervised manner [87]. GANs often suffer from mode collapse, where the model can only generate a single or small set of outputs. Wasserstein GAN improves the training by utilizing Wasserstein loss function, which measures the difference between the real and synthesized data distribution [88]. ProGAN tackles the training instability of GAN by progressively growing the generator and discriminator. The idea is that the model is scaled up gradually, starting with the simplest form of the problem, and little by little the problem’s complexity is increased as the training progresses [89]. StyleGAN leverages the progressive GAN’s approach and neural style transfer to improve the quality of the generated data [90]. The model is characterized by the independent manipulation of both style and content, allowing it to generate diverse styles and high-quality data.
4 State-of-the-art Deep Learning Applications
@@ 4.1 Computer Vision
Computer vision is an essential field in artificial intelligence (AI). It is a field of study that focuses on enabling computers to acquire, analyze and interpret visual inputs to derive meaningful information. The visual inputs can take many forms such as digital images, sequence of images or video and point cloud, and the source of these inputs can be camera, LiDAR and medical scanning machine. Deep learning, specifically CNN models have been widely used in real-world computer vision applications including image classification, object detection and image segmentation. This section discusses more details about the recent advancements in deep learning models that have been achieved over the past few years.
@@ 4.1.1 Image Classification
Image classification is a fundamental task in computer vision, which involves categorizing an image into one of predefined classes based on the visual content. The objective of image classification is to enable computers or machines to differentiate between objects within images, in a manner similar to how humans interpret visual information. Image classification is a crucial component in various applications such as robotics, manufacturing, and healthcare. LeNet-5, introduced in 1998, is one of the earliest convolutional neural networks that was successfully trained to classify handwritten digits. The model underwent a series of improvements, including the use of tanh and average pooling, which enhanced its ability to extract hierarchical features, ultimately improving overall performance. The model architecture comprises two convolutional layers, each with an average pooling layer, followed by two fully connected layers, including the output layer [96]. Since then, numerous CNN models have been proposed based on LeNet-5 for image classification [97, 98] but the most significant one is AlexNet in 2012, which saw a transformative breakthrough in deep learning. AlexNet is considered the first CNN model with a large number of parameters that significantly improved the performance of image classification on a very large dataset (ImageNet). The model won first place in ILSVRC 2012, improving the test error from the previous year by almost 10% [52]. Numerous significant CNN models have been introduced in subsequent ILSVRC competitions including ZFNet, VGG16, GoogleNet, ResNet and ResNext. In general, the research focused on increasing the number of layers, addressing the problem of vanishing gradient and diminishing of feature reuse.
@@
Vision transformers (ViT) offers an alternative to convolutional neural networks that have long been the dominant architecture for image classification, by leveraging self-attention mechanisms for scalable global representation learning. Despite its effectiveness, ViT is sensitive to hyperparameter optimization and substandard performance on smaller datasets [105]. Furthermore, ViT lacks the ability to leverage local spatial features which is inherent in convolutional neural networks [106]. Therefore, several studies attempt to incorporate convolutional layers into ViT architecture to improve its performance and robustness. In particular, conformer is a network architecture with two branches: CNN branch and a transformer branch to extract local and global features respectively [107]. Both branches are connected by two “bridges” of $1\times 1$ convolution and up or down sampling operations, allowing the branches to share their features and enhance the feature representation. Both branches output predictions which are combined to produce the final prediction.
A hybrid architecture, named MaxViT, combines convolutional networks and vision transformer to address the lack of scalability issues of self-attention mechanisms when the model is trained on large input size [108]. The improved vision transformer is composed of two modules whereby the first module attends local features in non-overlapping image patches and the global features are attended by processing a grid of sparse and uniform pixels. The transformer is stacked with a block of convolutional layers to extract local spatial features. The architecture of MaxViT is shown in Figure 22. Another study proposes a convolutional transformer network, introducing the depthwise convolutional block into the ViT [109]. This configuration allows the model to exploit the ability of convolutional networks to extract local spatial features while the ViT attends the extracted local features to focus on relevant information, enhancing the model’s ability to capture complex patterns and relationships. Specifically, instead of linear mapping, depth-wise convolutional mapping is used to generate query, key, and value matrices. Table 2 lists the summary of state-of-the-art image classification.
@@
Table 2: Summary of state-of-the-art image classification.
Ref. Description Results (Datasets and Metrics)
[99], 2022 The study introduced an encouraged loss that rewards well-classified examples with additive bonuses to enhance their contribution during training, addressing issues in representation learning, energy optimization, and makes the decision margin grows faster CIFAR-10
Accuracy: 92.97%
ImageNet
Accuracy: 76.43%
[100], 2023 The study introduced asymmetric polynomial loss to address class imbalance by decoupling gradient contributions from positive and negative instances, increasing the impact of updates for minority instances and helping the model focus on the more challenging, less frequent cases MS-COCO
mAP: 90.08
NUS-WIDE
mAP: 31.27
[101], 2023 The study introduced robust asymmetric loss to address class imbalance by emphasizing minority instances through an asymmetric weighting mechanism, allowing the model to focus more on the less frequent, yet critical minority instances ISIC2018
Accuracy: 0.852
APTOS2019
Accuracy: 0.826
[107], 2023 This study introduced dual network structure to extract local and global feature representations using convolutional layers and transformer ImageNet
Accuracy: 81.3 (Conformer-Ti)
Accuracy: 83.4 (Conformer-S)
Accuracy: 84.1 (Conformer-B)
[108], 2022 This study introduced MaxViT, a hybrid attention model, combining convolutional layers with transformer-based architecture whereby it uses convolutional layers in the early stages to capture local patterns efficiently, and then applies a multi-axis attention mechanism to capture both local and global dependencies more effectively ImageNet
Accuracy: 85.72 (MaxViT-T)
Accuracy: 86.19 (MaxViT-S)
Accuracy: 86.66 (MaxViT-B)
Accuracy: 86.70 (MaxViT-L)
[109], 2024 This study introduced convolutional transformer network that integrates convolutional layers and transformer encoder blocks whereby a transformer encoder block consists of multi-head self attention and feed-forward network layers with skip connections Diving 48
Accuracy: 77.3 (CTN-a), 78.2 (CTN-b)
Epic-Kitchens
Accuracy: 44.1 (CTN-a), 45.9 (CTN-b)
@@ 4.1.2 Object Detection
Deep learning plays a major role in significantly advancing the state-of-the-art in object detection performance. Region-based CNN (R-CNN) is the first breakthrough in object detection that combines CNN with selective region proposals [110]. The region proposals are the candidate bounding boxes serving as the potential region of interests (objects) within the input image, and the CNN are used to extract features from the region proposals and classify the regions for object detection. An improved model, named Fast R-CNN introduces two prediction branches: object classification and bounding box regression which improves the overall performance of object detection [111]. However, R-CNN and Fast R-CNN models are computationally expensive and slow, thus practically infeasible for real-time applications. Addressing this issue, Fast R-CNN is integrated with a region proposal network, referred to as Faster R-CNN [112]. The region proposal network (RPN) is used to efficiently generate region proposals for object detection. RPN takes an input image and output a set of rectangle object proposals, each with a confidence score to indicate the likelihood of an object’s presence. To this end, RPN introduces the concept of anchor boxes, whereby multiple bounding boxes of different aspect ratios are defined over the feature maps produced by the convolutional networks. These anchor boxes are then regressed over the feature maps to localize the objects, contributing to the improved speed and effectiveness of Faster R-CNN. The training of Faster R-CNN is divided into two stages. First, the RPN is pre-trained to generate the region proposals and then, the Fast R-CNN is trained using the region proposals generated by the RPN for object detection. The backbone network responsible for extracting the features for Faster R-CNN is either ZFNet or VGG16.
@@
In two-stage object detectors, the region proposals are generated first, and then used for object detection. The two-stage process is computationally intensive and infeasible for real-time object detection applications. You Only Look Once or YOLO proposes a one-stage detection by directly predicting bounding boxes and object’s confidence score in a single forward pass through the neural network [113]. This single pass architecture significantly reduces the computational complexity, making YOLO suitable for real-time object detection applications. In YOLO, the input image is divided into $S\texttimes S$ grids, each grid cell is responsible for detecting the objects present in the cell. Specifically, each grid cell predicts multiple bounding boxes and associated object’s confidence score, enabling simultaneous object detection across the entire image. Subsequent enhancements such as YOLOv3 [114] and YOLOv4 [115] are proposed, improving the model’s capability and accuracy. Single Shot Multibox Detector (SSD) is another one-stage detector, which aims to address the issue of real-time object detection [116]. SSD also eliminates the region proposal generation and directly predicts bounding boxes and confidence scores, reducing the computational complexity. To improve the overall performance, SSD produces the predictions from different levels of feature maps, allowing detection of objects of different sizes in the input image.
@@
A common issue in object detection problems is the extremely imbalanced ratio of foreground to background classes. Addressing this issue, RetinaNet introduces a loss function that is based on the cross entropy called focal loss. Focal loss reduces the loss contribution of easily classified objects, allowing the model training to focus on the difficult objects [117]. RetinaNet adopts the Feature Pyramid Network (FPN) [118] with ResNet as the backbone network for extracting the feature maps. FPN is a network architecture with a pyramid structure that efficiently captures multiscale feature representation, facilitating object detection across various sizes. To further improve the overall performance, EfficientDet introduces bidirectional FPN, which incorporates multi-level feature fusion to better capture multiscale feature representation [119]. Also, the model utilizes EfficientNet [120] as the backbone network to achieve a balance between computational efficiency and accuracy.
@@
Object detection performance often relies on a post-processing step called non-maximum suppression (NMS) to eliminate duplicate detections and select the most relevant bounding boxes. Specifically, NMS sorts all detection boxes based on their confidence scores, selects a box with the maximum score and discards the other boxes with a significant overlap with the selected box. This process is repeated on the remaining detection boxes. However, due to the inconsistency between the confidence score and the quality of object localization, NMS retains poorly localized bounding boxes with high confidence score while discarding more accurate predictions with poor confidence score. To mitigate this limitation, instead of discarding the neighboring boxes with significant overlap, soft-NMS applies Gaussian function to lower their confidence scores [121]. The idea is not to discard the neighboring bounding boxes, but gradually decline their scores based on the extent of the overlap with the selected box. This results in a smoother suppression, preserving the better-localized bounding boxes.
Adaptive NMS introduces an adaptive threshold for the suppression of bounding boxes [122]. The algorithm dynamically adjusts the threshold based on the level of overlapping of the selected box with the other bounding boxes. Similar method is reported in [123] whereby an adaptive NMS is proposed by dynamically adjusting the suppression criteria based on intersection over union (IoU) values. The IoU values are compared with a threshold and applies an additional iteration to vote for each detected proposal, ensuring better distinction between closely positioned objects. The method helps preserve multiple high-confidence bounding boxes, mitigating the issue of incorrectly merging adjacent objects into one. Redundant bounding boxes are often not filtered out due to their low IoU values with the best bounding box, leading to high false positives. In [124], a learning-based NMS is proposed to reduce false positives by integrating the NMS process into the model’s learning framework. Specifically, a novel NMS-aware loss is formulated that incorporates IoU to adjust bounding box weights, boosting negative attention for low IoU bounding boxes and enhancing positive weights for high IoU bounding boxes. Furthermore, regression assisted classification branch is introduced to aid classification by leveraging regression prediction relationships between bounding boxes and their best counterparts. In [125], the authors formulated similarity distance metric is proposed to evaluate the similarity between bounding boxes. The proposed method considers the location and shape of the boxes, and adapts dynamically to different datasets and object sizes. This method improves label assignment by learning hyperparameters automatically, eliminating the need for manual tuning and enhancing detection performance for tiny objects detection.
@@
Detection Transformer (DETR) is an end-to-end trainable object detection model that leverages the transformer architecture to eliminate the need for handcrafted components such as anchor boxes and non-maximum suppression [126]. The self-attention mechanism of the transformer captures the global context and relationships between different parts of the image, allowing it to localize the objects and remove duplicate predictions. The model is trained with a set of loss functions that perform bipartite matching between the predicted and ground truth objects. DETR uses ResNet as backbone network. Despite the success of DETR in simplifying and improving object detection tasks, DeTR suffers from a long training time and low performance at detecting small objects due to its reliance on the self-attention mechanism of the transformer, which lacks a multiscale feature representation. To mitigate this limitation, Deformable DETR introduces a multiscale deformable attention module which can effectively capture feature representation at different scales [127]. Furthermore, the attention module leverages deformable convolution, allowing the model to adapt to spatial variation and capture more informative features in the input data.
Dynamic DETR addresses the same issues by utilizing a deformable convolution-based FPN to learn multiscale feature representation [128]. The model replaces the transformer encoder with a convolution-based encoder to attend to various spatial features and channels. Moreover, an ROI-based dynamic attention is introduced at the transformer decoder, allowing the model to focus on the region of interests. This modification allows the model to effectively detect small objects and converge faster during training. The architecture of dynamic DETR is shown in Figure 23. DETR utilizes bipartite matching between the ground truth and the predicted objects to assign each ground truth object to a unique prediction, limiting the localization supervision. In [129], a hybrid matching scheme is introduced, combining the original one-to-one matching with an auxiliary one-to-many matching during training, improving training efficiency and detection accuracy without adding inference complexity. A training scheme known as Teach-DeTR is proposed to improve the overall performance of DeTR [130]. The training scheme leverages the predicted bounding boxes by other object detection models during the training by calculating the loss of one-to-one matching between the object queries and the predicted boxes. Table 3 lists the summary of state-of-the-art object detection.
@@
Table 3: Summary of state-of-the-art object detection.
Ref. Description Results (Datasets and Metrics)
[123], 2023 This study introduced an adaptive NMS method that dynamically adjusts suppression criteria based on intersection over union (IoU) values, applying an additional iteration to improve the distinction between closely positioned objects and preserve high-confidence bounding boxes, reducing false positives Rumex weeds
IoU: 91.2, mAP: 90
[124], 2024 This study proposed a learning-based NMS that integrates the NMS process into the model’s learning framework through a novel NMS-aware loss function, which adjusts bounding box weights using IoU to reduce false positives, while introducing a regression-assisted classification branch to enhance classification by leveraging bounding box prediction relationships CrowdHuman
mAP: 90.1
[125], 2024 This study proposed a similarity distance metric to evaluate bounding box similarity by considering location and shape, dynamically adapting to different datasets and object sizes, automatically learning hyperparameters to improve label assignment, and enhancing detection performance for tiny objects AI-TOD
mAP: 26.5, AP50: 57.7, AP75: 20.5
SODA-D
mAP: 32.8, AP50: 59.4, AP75: 31.3
VisDrone2019
mAP: 28.7, AP50: 50.3
[128], 2021 This study introduced Dynamic DETR, which enhances the original DETR by introducing dynamic attention mechanisms in both the encoder and decoder stages. This approach addresses limitations related to small feature resolution and slow training convergence, resulting in improved performance and efficiency in object detection tasks COCO2017
mAP: 49.3, AP50: 68.4, AP75: 53.6
[129], 2023 This study proposed a method that combines a one-to-one matching branch with an auxiliary one-to-many matching branch during training to improve the detection accuracy of DETR COCO2017
mAP: 59.4, AP50: 77.8, AP75: 65.4
[130], 2023 This study proposed a training scheme to enhances DETR by leveraging predicted bounding boxes from other teacher models including RCNN-based and DETR-based detectors using knowledge distillation. The method does not introduce extra parameters or computational overhead during inference COCO2017
mAP: 58.5, AP50: 77.4, AP75: 64.8